# Refined Risk Bounds for Unbounded Losses via Transductive Priors

### v1

@@ Appendix C Revisiting Existing Bounds for Sequential Linear Regression
$=\inf\limits_{\theta\in\mathbb{R}^{d}}\left(\sum_{t=1}^{T}(y_{t}-\langle x_{t},\theta\rangle)^{2}+\lambda\|\theta\|^{2}\right)+\sum\limits_{t=1}^{T}(h_{x_{t}}^{2}-h_{x_{t}})(\langle x_{t},\widehat{\theta}_{t}\rangle)^{2}+\sum\limits_{t=1}^{T}h_{x_{t}}y_{t}^{2}.$
@@
since $h_{x_{t}}\in[0,1]$. It is only left to bound $\sum\nolimits_{t=1}^{T}h_{x_{t}}y_{t}^{2}$, which, using (51) and $x\leq\log(1/(1-x))$ for $x\in(0,1)$, can be done as follows:
@@
In the following example, we demonstrate that Proposition C.1 implies the bound of Kivinen and Warmuth (1999) (see also (Forster, 1999)) for clipped ridge regression. This bound is also revisited in (Vovk, 2001, Theorem 4) using exp-concavity arguments applied to the squared loss. However, there is an issue with that proof. The exp-concavity argument in (Vovk, 2001, Theorem 4) necessitates that both $|y_{t}|\leq m$ and $|\langle x_{t},\theta\rangle|\leq m$ for $\theta$-s within the considered range. The second assumption is not mentioned in the statement of Theorem 4 in (Vovk, 2001), and indeed, it can be bypassed. Unlike the previous analysis, our approach does not rely on exp-concavity and, therefore, successfully recovers the claimed result without additional boundedness assumptions. The proof below was also presented in (Zhdanov and Vovk, 2009, Corollary 1), and we include it here primarily for the reader’s convenience.
Corollary C.2 (The regret of the clipped ridge predictor of Kivinen and Warmuth (1999)).
In the setup of Corollary C.1, we have
$\sum\limits_{t=1}^{T}\left(y_{t}-\operatorname{clip_{m}}\left(\langle x_{t},\widehat{\theta}_{t}\rangle\right)\right)^{2}\leq\inf\limits_{\theta\in\mathbb{R}^{d}}\left\{\sum_{t=1}^{T}(y_{t}-\langle x_{t},\theta\rangle)^{2}+\lambda\|\theta\|^{2}\right\}+4m^{2}\log\left(\det\left(I_{d}+\frac{1}{\lambda}\sum\limits_{t=1}^{T}x_{t}x_{t}^{\top}\right)\right).$
@@
Proof. Using Proposition C.1, the definition of the clipping operator and (53), we have
@@
Finally, to recover Theorem 2.1, originally due to Gaillard, Gerchinovitz, Huard, and Stoltz (2019), using Proposition C.1. We exactly repeat the lines of the proof of Corollary C.1 with the only change. Now, the matrix $A$ in Proposition C.1 is not equal to $\lambda I_{d}$, but is equal to $\lambda\sum\nolimits_{t=1}^{T}x_{t}x_{t}^{\top}$, which is assumed to be invertible. Finally, using an analog of (53), the total loss of the predictor is bounded by
@@
Similarly, repeating these lines but instead using the matrix $A=\lambda\sum\nolimits_{t=1}^{T}x_{t}x_{t}^{\top}$ in the proof of Corollary C.2, we obtain the following result.
Corollary C.3.
@@ Appendix D Achieving Similar Regret Bounds Using the $\varepsilon$-nets
Proof of Lemma D.1 Let $h(u)=-\log(1/(2T)+(1-1/T)u)$ for $u\in[0,1]$. Then
@@
Lemma D.2 (Lemma 6 of Drmota et al. (2024b)).
@@
One may alternatively prove Lemma D.2 using the bound on the covering numbers of VC-subgraph classes (see Györfi et al. (2002, Theorem 9.4)).
Lemma D.3.
@@
Proof of Lemma D.3 For any $\varepsilon>0$, let $\varepsilon^{\prime}=\varepsilon/(2T)$. Let $\mathcal{G}_{X,\varepsilon^{\prime}}$ be an $\varepsilon^{\prime}$-covering of the function class $\mathcal{G}_{X}$ in $L^{\infty}$ norm with at most $O((2T/\varepsilon)^{2d})$ elements. Then consider
@@
Then, for any $f_{\theta}\in\mathcal{F}_{X}$, there exists $\sigma(\langle\cdot,\widetilde{\theta}\rangle)\in\mathcal{G}_{X,\varepsilon}$ such that $\|\sigma(\langle\cdot,\theta\rangle)-\sigma(\langle\cdot,\widetilde{\theta}\rangle)\|_{\infty}\leq\varepsilon$. By Lemma D.1, we note that for any $x$,

### v3

@@ Appendix C Revisiting Existing Bounds for Sequential Linear Regression
$=\inf\limits_{\theta\in\mathbb{R}^{d}}\left(\sum_{t=1}^{T}(y_{t}-\langle x_{t},\theta\rangle)^{2}+\lambda\|\theta\|^{2}\right)+\sum\limits_{t=1}^{T}(h_{x_{t}}^{2}-h_{x_{t}})(\langle x_{t},\widehat{\theta}_{t}\rangle)^{2}$
$\qquad+\sum\limits_{t=1}^{T}h_{x_{t}}y_{t}^{2}.$
@@
since $h_{x_{t}}\in[0,1]$. It is only left to bound $\sum\nolimits_{t=1}^{T}h_{x_{t}}y_{t}^{2}$, which, using (C) and $x\leq\log(1/(1-x))$ for $x\in(0,1)$, can be done as follows:
@@
In the following example, we demonstrate that Proposition 36 implies the bound of Kivinen and Warmuth (1999) (see also (Forster, 1999)) for clipped ridge regression. This bound is also revisited in (Vovk, 2001, Theorem 4) using exp-concavity arguments applied to the squared loss. However, there is an issue with that proof. The exp-concavity argument in (Vovk, 2001, Theorem 4) necessitates that both $|y_{t}|\leq m$ and $|\langle x_{t},\theta\rangle|\leq m$ for $\theta$-s within the considered range. The second assumption is not mentioned in the statement of Theorem 4 in (Vovk, 2001), and indeed, it can be bypassed. Unlike the previous analysis, our approach does not rely on exp-concavity and, therefore, successfully recovers the claimed result without additional boundedness assumptions. The proof below was also presented in (Zhdanov and Vovk, 2009, Corollary 1), and we include it here primarily for the reader’s convenience.
Corollary 40 (The regret of the clipped ridge predictor of Kivinen and Warmuth (1999)).
In the setup of Corollary 39, we have
$\sum\limits_{t=1}^{T}\left(y_{t}-\operatorname{clip_{m}}\left(\langle x_{t},\widehat{\theta}_{t}\rangle\right)\right)^{2}$
$\qquad\leq\inf\limits_{\theta\in\mathbb{R}^{d}}\left\{\sum_{t=1}^{T}(y_{t}-\langle x_{t},\theta\rangle)^{2}+\lambda\|\theta\|^{2}\right\}+4m^{2}\log\left(\det\left(I_{d}+\frac{1}{\lambda}\sum\limits_{t=1}^{T}x_{t}x_{t}^{\top}\right)\right).$
@@
Proof. Using Proposition 36, the definition of the clipping operator and (55), we have
@@
Finally, to recover Theorem 1, originally due to Gaillard, Gerchinovitz, Huard, and Stoltz (2019), using Proposition 36. We exactly repeat the lines of the proof of Corollary 39 with the only change. Now, the matrix $A$ in Proposition 36 is not equal to $\lambda I_{d}$, but is equal to $\lambda\sum\nolimits_{t=1}^{T}x_{t}x_{t}^{\top}$, which is assumed to be invertible. Finally, using an analog of (55), the total loss of the predictor is bounded by
@@
Similarly, repeating these lines but instead using the matrix $A=\lambda\sum\nolimits_{t=1}^{T}x_{t}x_{t}^{\top}$ in the proof of Corollary 40, we obtain the following result.
Corollary 41.
@@ Appendix D Achieving Similar Regret Bounds Using the $\varepsilon$-nets
Proof of Lemma 42 Let $h(u)=-\log(1/(2T)+(1-1/T)u)$ for $u\in[0,1]$. Then
@@
Lemma 43 (Lemma 6 of Drmota et al. (2024b)).
@@
One may alternatively prove Lemma 43 using the bound on the covering numbers of VC-subgraph classes (see Györfi et al. (2002, Theorem 9.4)).
Lemma 44.
@@
Proof of Lemma 44 For any $\varepsilon>0$, let $\varepsilon^{\prime}=\varepsilon/(2T)$. Let $\mathcal{G}_{X,\varepsilon^{\prime}}$ be an $\varepsilon^{\prime}$-covering of the function class $\mathcal{G}_{X}$ in $L^{\infty}$ norm with at most $O((2T/\varepsilon)^{2d})$ elements. Then consider
@@
Then, for any $f_{\theta}\in\mathcal{F}_{X}$, there exists $\sigma(\langle\cdot,\widetilde{\theta}\rangle)\in\mathcal{G}_{X,\varepsilon}$ such that $\|\sigma(\langle\cdot,\theta\rangle)-\sigma(\langle\cdot,\widetilde{\theta}\rangle)\|_{\infty}\leq\varepsilon$. By Lemma 42, we note that for any $x$,